# CAMPBELL & REMIRO-AZÓCAR

### v1

@@ 3.3 Entropy balancing (matching-adjusted indirect comparison)
In addition, the entropy balancing estimator has a “linear double robustness” property as it is consistent under two distinct underlying data-generating models.40, 42, 43 Namely, it enables consistent estimation as long as either the true log-odds of the propensity score are linear across the specified balance functions (i.e., the implied propensity model is correctly specified); or, the true potential outcome under the active intervention is linear across the specified balance functions. Respectively, either $\operatorname{logit}(e_{i})=\alpha_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\alpha}$ for some parameters $\alpha_{0}$ and $\mathbf{\alpha}$; or, $\textrm{E}(Y^{1}_{i}\mid\mathbf{X}_{i})=\beta_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\beta}$ for some parameters $\beta_{0}$ and $\mathbf{\beta}$.
@@
For continuous covariates, it is common practice to only balance first-order moments (means) and, potentially, second-order sample moments (variances).25, 28, 38, 73 Covariate means are balanced by setting $\mathbf{c}(\mathbf{X}_{i})=c_{1}(\mathbf{X}_{i})=\mathbf{X}_{i}$ and $\hat{\mathbf{\theta}}=\hat{\mathbf{\theta}}_{1}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{X}_{i}$, in which case consistency is guaranteed as long as either the true propensity score model is $\operatorname{logit}(e_{i})=\alpha_{0}+\mathbf{X}_{i}^{\top}\mathbf{\alpha}$ for some parameters $\alpha_{0}$ and $\mathbf{\alpha}$; or, the potential outcome for the active intervention varies linearly with the covariates $\mathbf{X}_{i}$ that are mean-balanced: $\textrm{E}(Y^{1}_{i}\mid\mathbf{X}_{i})=\beta_{0}+\mathbf{X}_{i}^{\top}\mathbf{\beta}$ for some parameters $\beta_{0}$ and $\mathbf{\beta}$.
Both covariate means and variances are balanced by setting $\mathbf{c}(\mathbf{X}_{i})=[c_{1}(\mathbf{X}_{i}),c_{2}(\mathbf{X}_{i})]^{\top}=[\mathbf{X}_{i},\mathbf{X}^{2}_{i}]^{\top}$, $\hat{\mathbf{\theta}}=(\hat{\mathbf{\theta}}_{1},\hat{\mathbf{\theta}}_{2})^{\top}$, $\hat{\mathbf{\theta}}_{1}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{X}_{i}$ and $\hat{\mathbf{\theta}}_{2}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{X}_{i}^{2}$, in which case consistency is achieved if either the true propensity score model is $\textrm{logit}(e_{i})=\alpha_{0}+\mathbf{X}_{i}^{\top}\mathbf{\alpha_{1}}+(\mathbf{X}_{i}^{2})^{\top}\mathbf{\alpha_{2}}$ for some parameters $\alpha_{0}$, $\mathbf{\alpha_{1}}$ and $\mathbf{\alpha_{2}}$; or, the potential outcome for the active intervention varies quadratically with $\mathbf{X}_{i}$, such that $\textrm{E}(Y^{1}_{i}\mid\mathbf{X}_{i})=\beta_{0}+\mathbf{X}_{i}^{\top}\mathbf{\beta}_{1}+(\mathbf{X}_{i}^{2})^{\top}\mathbf{\beta}_{2}$ for some parameters $\beta_{0}$, $\mathbf{\beta_{1}}$ and $\mathbf{\beta_{2}}$. Balancing the means of the covariates and the squared covariates enforces that variances are balanced because $\textrm{Var}(X)=\textrm{E}(X^{2})-\textrm{E}(X)^{2}$.
3.4 Doubly robust augmented weighting estimators
@@
Unfortunately, the aforementioned balancing strategies are insufficient for consistency where the underlying data-generating models are more complex. One can consider various other non-linear transformations of $\mathbf{X}_{i}$, e.g., higher-order polynomial terms and flexible basis functions such as splines, and balance on the means of the transformed covariates. This would enable one to conjecture more flexible implicit data-generating models that are non-linear in the covariates. Moreover, one could go beyond the marginal moments of individual covariates and also balance joint covariate distributions; for instance, by including quantiles of interactions for pairs or triples of covariates.78 However, pursuing these more ambitious balancing strategies is often infeasible:
@@
  • As the number of balancing conditions increases, it is more likely that $\mathbf{\theta}$ falls outside the convex hull of $\mathbf{c}(\mathbf{X}_{i})$ for $i=1,\dots,n_{1}$.60 This implies that a feasible weighting solution to the convex optimization problem does not exist: there is no set of positive weights that can enforce balance in the required distributional features and the numerical optimization algorithm will fail to converge.66
@@
These difficulties motivate the explicit augmentation of the weighting estimators, as proposed by Funk et al (2011),79 allowing for a less restrictive outcome model that permits non-linear link functions and/or outcomes that depend on non-linear covariate transformations.
To construct such augmented estimators, we first postulate a model $m\left(\mathbf{X}_{i};\mathbf{\beta}\right)$ for the potential outcome expectation under the active intervention, conditional on covariates $\mathbf{X}_{i}$:
@@
where $q(\cdot)$ denotes an appropriate link function and $\mathbf{\beta}$ is a vector of model parameters encoding the covariate-outcome relationships. For instance, a logistic regression could be used for binary outcomes, such that the model is binomial, the link function is the logit and the potential outcome expectation is on the probability scale. We have assumed that the model for the conditional outcome expectation is parametric, but this need not necessarily be the case.80
The model in Equation 16 is fitted to the SAT participants $i=1,\dots,n_{1}$ using maximum-likelihood estimation, such that the fitted model $m\left(\mathbf{X}_{i};\hat{\mathbf{\beta}}\right)$ with parameter estimates $\hat{\mathbf{\beta}}$ is an estimator of the (transformed) conditional expectation $q\left(\textrm{E}(Y_{i}^{1}\mid\mathbf{X}_{i},S_{i}=1)\right)$. Based on the model, potential outcomes under the active intervention are predicted for each subject in the external control:
$\hat{Y}_{i}^{1}=q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})\right).$ (17)
@@
By averaging or “marginalizing” the potential outcome predictions generated by Equation 17 over the empirical covariate distribution of the external control arm, one obtains the “G-computation”, “regression standardization” or “plug-in G-formula” estimator33, 34, 44, 45, 46 for the mean potential outcome $\mu_{0}^{1}$ had subjects in the external control received the active intervention:
@@
The G-computation estimator for the ATC contrasts the average of potential counterfactual outcomes under the active intervention with the average of observed outcomes for the external control. We convert mean absolute outcomes to the additive scale imposed by link function $g(\cdot)$, prior to taking the difference between treatments on such scale:
@@
Note that the link function $g(\cdot)$ used for summarizing the treatment effect does not necessarily need to match the link function $q(\cdot)$ used for modeling.81
The G-computation estimators in Equations 18 and 19 rely on the outcome model in Equation 16 being correctly specified, in order to be consistent for the mean absolute outcome ${\mu}_{0}^{1}$ and the ATC, respectively. Assuming correct model specification for all estimators, G-computation is more precise and efficient than weighting, particularly when poor overlap leads to large reductions in effective sample size.34, 46 However, the increase in precision is achieved by implicit extrapolation into non-overlapping regions of the covariate space, hiding underlying failures of the positivity assumption.48 Model misspecification bias is almost impossible to diagnose in extrapolated regions, and there is typically no inflation of the variance to reflect the extrapolation uncertainty. Our proposed augmented weighting estimators will not use the outcome model to extrapolate, but to gain bias-robustness – and, potentially, some precision35 – with respect to their weighting counterparts.
Firstly, based on Funk et al (2011) and Shinozaki and Matsuyama (2015),79, 82 we propose combining the modeling approach to inverse odds weighting, described in Section 3.2, with G-computation. Specifically, suppose we have fitted the outcome model in Equation 16 to the SAT. We now let $q^{-1}\left(m(\mathbf{X}_{i};\hat{\mathbf{\beta}})\right)$ be a prediction of potential outcome $Y_{i}^{1}$ for the active intervention based on the fitted regression, not only for the external control subjects but for all subjects $i=1,\dots,n$ in the SAT and the external control. The propensity score weights $\hat{w}_{i}$ derived for $i=1,\dots,n_{1}$ are used to add an error-correcting term to the G-computation estimator. The resulting augmented inverse odds weighting estimator for the mean absolute outcome $\mu_{0}^{1}$ is:
@@
where $\epsilon_{i}^{1}=Y_{i}-\hat{Y}_{i}^{1}$ is a residual term for subject $i=1,\dots,n_{1}$ in the SAT. Note that this estimator exactly corresponds to the doubly robust estimator proposed by Shinozaki and Matsuyama (2015),82 except that their estimator is for $\textrm{E}(Y^{0}|S=1)$ while the estimator in Equation 20 is for $\textrm{E}(Y^{1}|S=0)$ (the estimand ultimately targeted by Shinozaki and Matsuyama is the ATT but ours is the ATC).
In Equation 20, the G-computation estimator has been augmented with a weighted average of the residuals for the SAT subjects. Such a weighted average is a “one-step” correction term for the potential bias of the G-computation estimator.83 The corresponding bias-corrected estimator for the ATC is:
@@
on the additive scale imposed by link function $g(\cdot)$. Normalizing the weights so that they sum to one and to ensure bounded estimates, we obtain the alternative estimators:
@@
$\widehat{\textrm{ATC}}=g\underbrace{\left(\frac{\sum_{i=1}^{n_{1}}\hat{w}_{i}\epsilon_{i}^{1}}{\sum_{i=1}^{n_{1}}\hat{w}_{i}}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (23)
Equations 22 and 23 should provide improved finite sample properties and more stable and precise estimation than Equations 20 and 21, respectively.
Our novel contribution is combining the entropy balancing-based MAIC approach, described in Section 3.3, with the G-computation estimator. Again, based on the outcome model fitted to the SAT, let $q^{-1}\left(m(\mathbf{X}_{i};\hat{\mathbf{\beta}})\right)$ be a prediction of the potential outcome $Y_{i}^{1}$ under the active intervention for all subjects $i=1,\dots,n$ in the SAT and the external control. In this case, it is the MAIC weights, $\hat{\omega}_{i}$ for $i=1,\dots,n_{1}$, derived in Equation 14, that are used to add an error-correcting term to the G-computation estimator. The resulting augmented MAIC estimator of the mean absolute outcome $\mu_{0}^{1}$ is:
$\hat{\mu}_{0}^{1}=$ $\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\left(Y_{i}-\hat{Y}^{1}_{i}\right)+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}$
$=$ $\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\epsilon_{i}^{1}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1},$ (24)
@@
The corresponding bias-corrected augmented MAIC estimator for the ATC is:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\epsilon_{i}^{1}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (25)
@@
on the additive scale imposed by link function $g(\cdot)$. We conjecture that the augmented MAIC estimators in Equations 24 and 25 will perform better statistically than the augmented estimators based on the modeling approach to weighting, which could exhibit erratic performance with highly variable weights, particularly if these are combined with a misspecified outcome model.84, 85 We expect the augmented MAIC estimators to inherit the more attractive properties of the entropy balancing weights: (1) lower susceptibility to bias by directly enforcing covariate balance; and (2) greater stability, translating into larger effective sample sizes after weighting and enhanced precision in estimation.
@@
The augmented weighting estimators in Equations 20, 22 and 24 are doubly robust for the mean potential outcome $\mu^{1}_{0}$. That is, they estimate $\mu^{1}_{0}$ consistently as long as either the propensity score model for data source assignment or the outcome model is correctly specified, but not necessarily both. By extension, the augmented weighting estimators in Equations 21, 23 and 25 are doubly robust for the ATC, under the assumption that $\hat{\mu}_{0}^{0}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}$ is consistent for $\mu_{0}^{0}$. We provide an intuitive heuristic to demonstrate double robustness.
Consider the simple scenario where we have a binary outcome, $Y$, and a single discrete covariate, $X$, and the augmented weighting estimator in Equation 24 is re-written as:
$\hat{\mu}_{0}^{1}=\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\left(Y_{i}-\hat{Y}^{1}_{i}\right)+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}.$ (26)
@@
If the outcome model is correctly specified, the expectation of the first summation in Equation 26 converges to zero as $n_{1}\rightarrow\infty$ because $\hat{Y}_{i}^{1}\rightarrow Y_{i}$ and the terms inside the summation cancel out, irrespective of any postulated propensity score model. The second summation is equivalent to the G-computation estimator and is consistent for $\mu_{0}^{1}$ because the outcome model is correct. Consequently, $\hat{\mu}_{0}^{1}\rightarrow\mu_{0}^{1}$, and $\rightarrow\textrm{ATC}$ ATC (assuming $\hat{\mu}_{0}^{0}\rightarrow\mu_{0}^{0}$).
Conversely, if the propensity score model is correctly specified but the outcome model is incorrect, the first summation consistently cancels out the bias produced by the G-computation estimator in the second summation and the remainder term is exactly equal to the non-augmented weighting estimator, which converges to $\mu_{0}^{1}$ as $n_{1}\rightarrow\infty$ because the propensity score model is correct. To illustrate this, consider rearranging Equation 26 to:
$\hat{\mu}_{0}^{1}$ $=\sum_{i=1}^{n_{1}}\hat{\omega}_{i}Y_{i}-\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\hat{Y}^{1}_{i}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}$
$=\underbrace{\sum_{i=1}^{n_{1}}\hat{\omega}_{i}Y_{i}}_{C}+\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}-\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\hat{Y}^{1}_{i}\right)}_{D}.$ (27)
First, $C$ is equivalent to the MAIC estimator and is consistent for $\mu_{0}^{1}$ because the propensity score model is correct. Expanding the summations in $D$ over $i=1,...,n$, we have:
$D$ $=\left(\sum_{i=1}^{n}\frac{(1-S_{i})\hat{Y}_{i}^{1}}{n_{0}}-\frac{S_{i}\hat{\nu}_{i}\hat{Y}_{i}^{1}}{n_{0}}\right)$ (28)
$=\frac{1}{n_{0}}\left(\sum_{i=1}^{n}{\hat{Y}_{i}^{1}}\left((1-S_{i})-S_{i}\hat{\nu}_{i}\right)\right),$ (29)
where $\hat{\nu}_{i}=n_{0}\hat{\omega}_{i}$, for $i$ in 1,…,$n$. Then, suppose that the outcome model for $Y^{1}$ is independent of $S$ and that $Y^{1*}_{i}$ is the large sample limit of $\hat{Y}^{1}_{i}$. Also, suppose that the large sample limit of $\hat{\nu}_{i}$ is $\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}$, since the propensity score model is correctly specified. Then:
$\textrm{E}\left(\hat{Y}_{i}^{1}\left((1-S_{i})-S_{i}\hat{\nu}_{i}\right)\right)$ $=\textrm{E}\left(\textrm{E}\left(\hat{Y}_{i}^{1}\left((1-S_{i})-S_{i}\hat{\nu}_{i}\right)\mid X\right)\right)$
$=\textrm{E}\left(\textrm{E}(\hat{Y}_{i}^{1}\mid X)\times\textrm{E}\left(\left((1-S_{i})-S_{i}\hat{\nu}_{i}\right)\mid X\right)\right)$
$\rightarrow\textrm{E}\left(Y^{1*}_{i}\sum_{x}{\textrm{Pr}(S=0|X=x)\textrm{Pr}(X=x)-\textrm{Pr}(S=1|X=x)\textrm{Pr}(X=x)\frac{\textrm{Pr}(S=0|X=x)}{\textrm{Pr}(S=1|X=x)}}\right)$
$=\textrm{E}\left(Y^{1*}_{i}\sum_{x}{\textrm{Pr}(S=0|X=x)\textrm{Pr}(X=x)-\textrm{Pr}(X=x)\textrm{Pr}(S=0|X=x)}\right)$
$=\textrm{E}\left(Y^{1*}_{i}\times 0\right)$
$=0.$ (30)
Consequently, $D\rightarrow 0$ which implies that $\hat{\mu}_{0}^{1}\rightarrow\mu_{0}^{1}$, and $\rightarrow\textrm{ATC}$ ATC having assumed $\hat{\mu}_{0}^{0}\rightarrow\mu_{0}^{0}$.
@@
While all the augmented weighting estimators described in this section are doubly robust, the augmented MAIC estimator defined in Equation 25 is arguably more robust to model misspecification bias because it is consistent under a greater number of distinct underlying data-generating mechanisms. Namely, the augmented MAIC estimator is consistent as long as either: (1) the log-odds of the propensity score are linear across the covariate balance functions; (2) the potential outcome under the active intervention is linear across the covariate balance functions; or (3) the explicit augmentation model for the potential outcome under the active intervention is correctly specified. The estimation of the weights is consistent as long as either the first or the second condition holds. Conversely, the augmented estimators based on the modeling approach to weighting (defined in Equations 20 to 23) are consistent as long as either the first or the third condition holds, with the first condition being necessary for consistent estimation of the weights.
3.5 Other augmented weighting estimators

### v2

@@ 3.3 Entropy balancing (matching-adjusted indirect comparison)
MAIC (entropy balancing) has a “linear double robustness” property as it is consistent under two distinct underlying data-generating models:josey2021transporting, cheng2023double, zhao2017entropy (1) the true log-odds of the propensity score are linear across the specified balance functions; or (2) the true potential outcome under the active intervention is linear across the specified balance functions. Respectively, either $\operatorname{logit}(e_{i})=\alpha_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\alpha}$ for some parameters $\alpha_{0}$ and $\mathbf{\alpha}$; or, $\textrm{E}(Y^{1}_{i}\mid\mathbf{X}_{i})=\beta_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\beta}$ for some parameters $\beta_{0}$ and $\mathbf{\beta}$.
@@
For continuous covariates, it is common practice to only balance first-order sample moments (means) by setting $\mathbf{c}(\mathbf{X}_{i})=c_{1}(\mathbf{X}_{i})=\mathbf{X}_{i}$ and $\hat{\mathbf{\theta}}=\hat{\mathbf{\theta}}_{1}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{X}_{i}$, in which case consistency is guaranteed as long as either the true propensity score model is $\operatorname{logit}(e_{i})=\alpha_{0}+\mathbf{X}_{i}^{\top}\mathbf{\alpha}$; or, the potential outcome for the active intervention varies linearly with the covariates $\mathbf{X}_{i}$ that are mean-balanced: $\textrm{E}(Y^{1}_{i}\mid\mathbf{X}_{i})=\beta_{0}+\mathbf{X}_{i}^{\top}\mathbf{\beta}$. An alternative strategy is to also balance second-order sample moments (variances) by setting $\mathbf{c}(\mathbf{X}_{i})=[c_{1}(\mathbf{X}_{i}),c_{2}(\mathbf{X}_{i})]^{\top}=[\mathbf{X}_{i},\mathbf{X}^{2}_{i}]^{\top}$, $\hat{\mathbf{\theta}}=(\hat{\mathbf{\theta}}_{1},\hat{\mathbf{\theta}}_{2})^{\top}$, $\hat{\mathbf{\theta}}_{1}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{X}_{i}$ and $\hat{\mathbf{\theta}}_{2}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{X}_{i}^{2}$, in which case consistency is achieved if either the true propensity score model is $\textrm{logit}(e_{i})=\alpha_{0}+\mathbf{X}_{i}^{\top}\mathbf{\alpha_{1}}+(\mathbf{X}_{i}^{2})^{\top}\mathbf{\alpha_{2}}$; or, the potential outcome for the active intervention varies quadratically with $\mathbf{X}_{i}$, such that $\textrm{E}(Y^{1}_{i}\mid\mathbf{X}_{i})=\beta_{0}+\mathbf{X}_{i}^{\top}\mathbf{\beta}_{1}+(\mathbf{X}_{i}^{2})^{\top}\mathbf{\beta}_{2}$. Balancing the means of the covariates and the squared covariates enforces that variances are balanced because $\textrm{Var}(X)=\textrm{E}(X^{2})-\textrm{E}(X)^{2}$.
@@
One could conjecture more flexible implicit data-generating models by considering other non-linear transformations of $\mathbf{X}_{i}$, e.g., higher-order polynomial terms and flexible basis functions such as splines, and balancing on the means of the transformed covariates. Moreover, one could go beyond the marginal moments of individual covariates and also balance joint covariate distributions; for instance, by including quantiles of interactions for pairs or triples of covariates.zubizarreta2015stable However, pursuing these more ambitious balancing strategies is often infeasible:
@@
  • As the number of balancing conditions increases, it is more likely that $\mathbf{\theta}$ falls outside the convex hull of $\mathbf{c}(\mathbf{X}_{i})$ for $i=1,\dots,n_{1}$.glimm2022geometric This implies that a feasible weighting solution to the convex optimization problem does not exist: there is no set of positive weights that can enforce balance in the required distributional features and the numerical optimization algorithm will fail to converge.chattopadhyay2020balancing
@@
3.4 G-computation
An alternative to weighting involves postulating a model $m\left(\mathbf{X}_{i};\mathbf{\beta}\right)$ for the potential outcome expectation under the active intervention, conditional on covariates $\mathbf{X}_{i}$:
@@ 3.4 G-computation
where $q(\cdot)$ denotes an appropriate link function and $\mathbf{\beta}$ is a vector of model parameters encoding the covariate-outcome relationships. For instance, a logistic regression could be used for binary outcomes, such that the model is binomial, the link function is the logit and the potential outcome expectation is on the probability scale. We have assumed that the model for the conditional outcome expectation is parametric, but this need not necessarily be the case.kern2016assessing
[…]
@@
By averaging (“marginalizing”) the potential outcome predictions generated by Equation 14 over the empirical covariate distribution of the external control, one obtains the G-computation estimator,loiseau2022external, wang2017g, robins1986new, keil2014parametric, remiro2022parametric also known as “regression standardization” or “plug-in G-formula”, for the mean potential outcome $\mu_{0}^{1}$ had subjects in the external control received the active intervention:
@@
Consequently, mean absolute outcomes are converted to the additive scale imposed by link function $g(\cdot)$, prior to taking the difference between treatments on such scale to derive the G-computation estimator for the ATC:
@@
Note that the link function $g(\cdot)$ used for summarizing the treatment effect does not necessarily need to match the link function $q(\cdot)$ used for modeling.remiro2024transportability
The readership may be more familiar with the well-known G-computation estimator for the ATE, which contrasts averages of predicted potential outcomes between both treatment groups.ren2023comparing Our G-computation estimator in Equation 16 contrasts the average of predicted counterfactual outcomes under the active intervention with the average of observed factual outcomes for the external control group. To estimate the ATC, only potential outcomes for the SAT subjects need to be predicted because all potential outcomes under the control have already been observed for the subjects in the external control.wang2017g
The G-computation estimators in Equations 15 and 16 rely on the outcome model in Equation 13 being correctly specified, in order to be consistent for the mean absolute outcome ${\mu}_{0}^{1}$ and the ATC, respectively. Assuming correct model specification for all estimators, G-computation is more precise and efficient than weighting, particularly when poor overlap leads to large reductions in effective sample size.loiseau2022external, remiro2022parametric However, the increase in precision is achieved by implicit extrapolation into non-overlapping regions of the covariate space, hiding underlying failures of the positivity assumption.vo2023cautionary Model misspecification bias is almost impossible to diagnose in extrapolated regions, and there is typically no inflation of the variance to reflect the extrapolation uncertainty.
3.5 Doubly robust augmented weighting estimators
The limitations of the methods in Section 3.2, Section 3.3 and Section 3.4 motivate the explicit augmentation of the weighting estimators,funk2011doubly allowing for a less restrictive outcome model that permits non-linear link functions and/or outcomes that depend on non-linear covariate transformations. Our proposed augmented weighting estimators will not use the outcome model to extrapolate, but to gain bias-robustness – and, potentially, some precisionlunceford2004stratification – with respect to their weighting counterparts. To construct such augmented estimators, we proceed as follows.
[…]
@@ 3.5 Doubly robust augmented weighting estimators
where $\epsilon_{i}^{1}=Y_{i}-\hat{Y}_{i}^{1}$ is a residual term for subject $i=1,\dots,n_{1}$ in the SAT. Note that this estimator exactly corresponds to the doubly robust estimator proposed by Shinozaki and Matsuyama (2015),shinozaki2015brief except that their estimator is for $\textrm{E}(Y^{0}|S=1)$ while the estimator in Equation 17 is for $\textrm{E}(Y^{1}|S=0)$. The estimand ultimately targeted by Shinozaki and Matsuyama is the ATT but ours is the ATC.
In Equation 17, the G-computation estimator has been augmented with a weighted average of the residuals for the SAT subjects. We shall explain later in this section why such a weighted average is an error-correcting term for the potential bias of the G-computation estimator.benkeser2017doubly The corresponding DR augmented IOW estimator for the ATC is:
@@
on the additive scale imposed by link function $g(\cdot)$. As for the non-augmented IOW estimators in Section 3.2, we can normalize or stabilize the weights so that they sum to one, thereby ensuring bounded estimates and improving the finite sample properties of the estimator.dahabreh2020extending Using the normalized weights, we obtain the DR augmented normalized IOW estimators:
@@
$\widehat{\textrm{ATC}}=g\underbrace{\left(\frac{\sum_{i=1}^{n_{1}}\hat{w}_{i}\epsilon_{i}^{1}}{\sum_{i=1}^{n_{1}}\hat{w}_{i}}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}}.$ (20)
Equations 19 and 20 should provide more stable and precise estimation than Equations 17 and 18, respectively.dahabreh2020extending
[…]
$\hat{\mu}_{0}^{1}=$ $\sum_{i=1}^{n_{1}}\hat{v}_{i}\left(Y_{i}-\hat{Y}^{1}_{i}\right)+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}$
$=$ $\sum_{i=1}^{n_{1}}\hat{v}_{i}\epsilon_{i}^{1}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1},$ (21)
@@
The corresponding DR augmented MAIC estimator for the ATC is:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\sum_{i=1}^{n_{1}}\hat{v}_{i}\epsilon_{i}^{1}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (22)
@@
on the additive scale imposed by link function $g(\cdot)$. We conjecture that the DR augmented MAIC estimators in Equations 21 and 22 will perform better statistically than the augmented estimators based on the modeling approach to weighting, which could exhibit erratic performance with highly variable weights, particularly if these are combined with a misspecified outcome model.robins2007comment, kang2007demystifying We expect the DR augmented MAIC estimators to inherit the more attractive properties of the entropy balancing weights: (1) lower susceptibility to bias by directly enforcing covariate balance; and (2) greater stability, translating into larger effective sample sizes after weighting and enhanced precision in estimation.
@@
The augmented weighting estimators in Equations 17, 19 and 21 are doubly robust for the mean potential outcome $\mu^{1}_{0}$. That is, they estimate $\mu^{1}_{0}$ consistently as long as either the propensity score model for data source assignment or the outcome model is correctly specified, but not necessarily both. Note that, contrary to the approach that we shall describe in Section 3.6, this double robustness does not depend on using a canonical link function for the outcome model. In the Appendix, we provide an intuitive heuristic to demonstrate double robustness, which illustrates why the weighted average of residuals is an error-correcting term for the potential bias of G-computation. The heuristic is summarized as follows. Consider that all the augmented weighting estimators have the general form:
$\hat{\mu}_{0}^{1}=\sum_{i=1}^{n_{1}}\hat{u}_{i}\left(Y_{i}-\hat{Y}^{1}_{i}\right)+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1},$ (23)
for a generic weight estimate $\hat{u}_{i}$, where $\hat{u}_{i}=\hat{w}_{i}/n_{0}$ for the DR augmented IOW estimator in Equation 17, $\hat{u}_{i}=\hat{w}_{i}/\sum_{i=1}^{n_{1}}\hat{w}_{i}$ for the DR augmented normalized IOW estimator in Equation 19, and $\hat{u}_{i}=\hat{v}_{i}$ for the DR augmented MAIC estimator in Equation 21.
@@
If the outcome model is correctly specified, the expectation of the first summation in Equation 23 converges to zero as $n_{1}\rightarrow\infty$ because $\hat{Y}_{i}^{1}\rightarrow Y_{i}$ and the terms inside the summation cancel out, irrespective of any postulated propensity score model for data source assignment. The second summation is equivalent to the G-computation estimator and is consistent for $\mu_{0}^{1}$ because the outcome model is correct. Consequently, $\hat{\mu}_{0}^{1}\rightarrow\mu_{0}^{1}$ and $\widehat{\textrm{ATC}}\rightarrow\textrm{ATC}$, under the assumption that $\hat{\mu}_{0}^{0}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}$ is consistent for $\mu_{0}^{0}$. Conversely, if the propensity score model is correctly specified but the outcome model is incorrect, the first summation consistently cancels out the bias produced by the G-computation estimator in the second summation and the remainder term is exactly equal to the non-augmented weighting estimator, which converges to $\mu_{0}^{1}$ as $n_{1}\rightarrow\infty$ because the propensity score model is correct. Similarly, $\widehat{\textrm{ATC}}\rightarrow\textrm{ATC}$ (assuming $\hat{\mu}_{0}^{0}\rightarrow\mu_{0}^{0}$). Hence, the augmented weighting estimators in Equations 18, 20 and 22 are doubly robust for the ATC.
@@
While all the augmented weighting estimators described in this section are doubly robust, the DR augmented MAIC estimators defined in Equations 21 and 22 are arguably more robust to model misspecification bias because they are consistent under a greater number of distinct underlying data-generating mechanisms. Namely, DR augmented MAIC is consistent as long as either: (1) the log-odds of the propensity score are linear across the covariate balance functions; (2) the potential outcome under the active intervention is linear across the covariate balance functions; or (3) the explicit augmentation model for the potential outcome under the active intervention is correctly specified. The estimation of the weights is consistent as long as either the first or the second condition holds. Conversely, the augmented estimators based on the modeling approach to weighting (defined in Equations 17 to 20) are consistent as long as either the first or the third condition holds, with the first condition being necessary for consistent estimation of the weights.
3.6 Weighted G-computation